# Estimating Causal Effects from Learned Causal Networks

### v1

@@ 1.  Initialize: $BIC_{\mathcal{B}}\leftarrow\infty$,
  4.     $(\mathcal{B}^{\prime},LL_{\mathcal{B}^{\prime}})\leftarrow\max\limits_{LL}\left\{\mathrm{EM}(\mathcal{G},\mathcal{D},k)|\textbf{for }i=1\text{ to }10\right\}$
@@ 7.          $\mathcal{B}\leftarrow\mathcal{B}^{\prime}$, $BIC_{\mathcal{B}}\leftarrow BIC_{\mathcal{B}^{\prime}}$
  10:  $\mathcal{B}_{{\boldsymbol{X}=\boldsymbol{x}}}$leftarrow generate truncated CBN from $\mathcal{B}$.
  11:  return   $\leftarrow$ evaluate $P_{\mathcal{B}_{\boldsymbol{X}=\boldsymbol{x}}}({\boldsymbol{Y}})$
@@
Since we use EM for learning we call our algorithm EM for Causal Inference (EM4CI), described in Algorithm 1. Its input is a causal diagram $\mathcal{G}$, samples $\mathcal{D}$ from $P(\boldsymbol{V})$, and a query $Q=P({\boldsymbol{Y}}\mid do({\boldsymbol{X}=\boldsymbol{x}}))$. After initialization, step 2 checks if the query is identifiable. This can be accomplished by any variant of the well-known ID algorithm [27, 24], or via the Do calculus [23]. If the query is not identifiable, the algorithm terminates with failure. Otherwise steps 3-9 provide the iterative learning scheme. Given samples $\mathcal{D}$ from $P(\boldsymbol{V})$, we use EM to learn a CBN $\mathcal{B}$ over the graph $\mathcal{G}$ with latent domain size $k$. EM outputs the parameters for a network, $\mathcal{B}^{\prime}$, and its corresponding log-likelihood, $LL_{\mathcal{B}^{\prime}}$, with which we calculate the $BIC_{\mathcal{B}^{\prime}}$ score. Since EM’s performance is known to be sensitive to initialization, we perform ten runs of EM starting from randomly generated initial parameters, and retain the model with the highest likelihood as the candidate for the current latent domain size $k$. We then increase $k$, and repeat the process. If the $BIC$ score decreases (step 6), we adopt the new candidate network and the process continues with increased $k$; otherwise the current $\mathcal{B}$ is selected as the final learned network. Step 10 generates the truncated CBN $\mathcal{B}_{\boldsymbol{X}=\boldsymbol{x}}$, and step 11 employs a standard PGM inference algorithm, like join-tree decomposition [6, 8], to the CBN $\mathcal{B}_{\boldsymbol{X}=\boldsymbol{x}}$ to compute $P_{\boldsymbol{X}=\boldsymbol{x}}({\boldsymbol{Y}})$.
@@
Determining identifiability is polynomial in the graph size [27]. The complexity of EM (steps 3-9) depends on the sample size $|D|$, the variables’ domain sizes, and the number of iterations needed for convergence which is hard to predict, but complexity-wise provide only a constant factor and can be ignored. The most extensive computation in each iteration of EM is the Expectation step, which requires probabilistic inference and can be accomplished in time and memory exponential in the treewidth of the graph. Otherwise, approximation algorithms such as belief propagation may be used. In our experiments, exact inference was always possible. Thus, each iteration of EM takes $O(|D|\cdot nl^{w})$, where $n=|V\cup U|$ is the number of variables, $l$ bounds the variable domain sizes, and $w$ is the treewidth. For $T$ iterations we find that the complexity of EM is $O(T\cdot|D|\cdot n\cdot l^{w})$. Step 11 of probabilistic inference is $O(n\cdot l^{w})$ if performing exact inference. In summary, the complexity of algorithm EM4CI is $O(T\cdot|D|\cdot n\cdot l^{w})$. Thus, the algorithm is most effective for models with bounded treewidth. Note that the cost of the EM learning process can be amortized over multiple queries.
@@
Table 1: Estimand Expressions for Models 1-8 in Figure 2(a) & Table 3.
@@ 4 Empirical Evaluation
We perform an extensive empirical evaluation of EM4CI and compare against estimand-based approaches such as the brute-force empirical plug-in method and a state-of-the-art scheme called weighted empirical risk minimization (WERM) [13]. All experiments were run on a 2.66 GHz processor with 8 GB of memory.
@@
Figure 3: Example larger synthetic models.
@@
We generate each triplet benchmark instance by first choosing a causal diagram and a query. Then, picking domain sizes of observed and latent variables, we generate the conditional probability tables (CPTs), one per variable and its parents, to yield the full CBN. Given the full model we generate samples from its observational distribution by forward sampling [6] over all variables, then discard the values of the latent variables. We compute the exact answer to the target query on the full CBN via an exact algorithm (e.g., join-tree [6, 8]).
The causal diagrams were selected in two ways. First, we take a set of 8 small diagrams from the causality literature [23, 14]; see Figure 2(a) and the Appendix. Second, we examine three scalable classes of graphs whose treewidths vary but can be controlled. These are: chain networks (treewidth 2; Figure 3(c)), diamond networks (treewidth 5; Figure 3(a)), and cone-cloud networks (treewidth $O(\sqrt{n})$, Figure 3(b)), abbreviated CH, D, CC respectively.
In the small diagrams (e.g., Figure 2(a)), we set the domain size of the observed variables to $d=2$, and of the latent variables to $k=10$. For the chain, diamond, and cone-cloud models, we use $(d,k)=(4,10)$. The parameters of each CPT were generated by sampling from a Dirichlet distribution [6] with parameters $\alpha=[\alpha_{1},\ldots,\alpha_{n}]$. We chose $\alpha_{i}\in[0,1)$ uniformly at random in order to generate conditional probabilities near the edges of the simplex (i.e., far from uniform).
@@
We also test on four networks created for real-world domains. The “Alarm” network was developed for on-line monitoring of patients in intensive care units [2]; the “A” network, from the UAI literature, is synthetic but known to be daunting to exact algorithms [17]; the “Barley” network was built for a decision support system for growing barley without pesticides [18]; and the “Win95” model is an expert system for printer troubleshooting in Windows 95 . Since no variables in these networks are specified as latent, we chose source vertices with at least two children to be latent confounding variables. We generated data by forward sampling and then discarding the latent variables’ values.
@@
We hand selected identifiable queries, aiming for non-trivial ones and prioritizing those that appear to be complex. The queries and their corresponding estimands are given in the Appendix. We selected several identifiable queries per model to better evaluate the methods, as well as to emphasize that the learning time can be amortized over multiple queries.
@@
We used the ID algorithm in the causal effect package [29] to compute the estimands. Subsequently, the plug-in method evaluates the estimand directly from the observational data, by plugging in the empirical conditional probabilities. Our plug-in implementation uses sparse table representations in Python, to ensure that the representation size of each estimated probability term is linear in the data size, rather than exponential in the number of variables. We also compare against a state-of-the-art scheme called WERM [13] using the author-provided implementation in R. A summary of the algorithms and implementations is given in Table 2.
@@
We report the results of EM4CI along the two phases of the algorithm; the learning phase (steps 3-9) done using the EM algorithm and the inference phase (steps 10-11) done using the join tree algorithm. For the learning process, we report the selected latent domain sizes and the total time at termination. For the inference phase we report the mean absolute deviation (mad) between the true answer and the estimated answer. The measure $mad$ for a query $P({Y}\mid do({\boldsymbol{X}}))$ is computed by averaging the absolute error over all single-value queries over all instantiations of the intervened and queried variables, $P({Y=y}\mid do({\boldsymbol{X}}=\boldsymbol{x}))$ for $\boldsymbol{x},y\in\mathcal{D}({\boldsymbol{X}})\times\mathcal{D}(Y)$. We also report average inference time over the same set of query instantiations. In our experiments we increased $k$ by twos ($k=2,4,6,8,10\ldots)$ in the EM4CI algorithm.
@@
1 2 0.0037 0.4759 0.0104 1.9 2 0.0032 3.1 0.0025 2.3
2 2 0.1832 1.8643 0.1436 2.3 2 0.0490 8.4 0.0867 2.0
3 2 0.1288 0.9288 0.0569 1.1 2 0.0040 3.6 0.0039 0.7
4 2 0.1819 1.8169 0.1469 2.3 2 0.1438 12.0 0.0704 2.1
5 2 0.4910 1.6539 0.5000 2.0 2 0.0044 17.3 0.0058 2.2
6 2 0.2663 0.3004 0.3930 2.0 2 0.1263 0.5 0.1319 2.1
7 2 0.2520 0.7757 0.2509 1.9 2 0.0891 7.1 0.0238 2.0
8 2 0.1372 0.6348 0.1579 2.0 2 0.2340 4.7 0.1303 1.9
Table 4: Results of absolute error on Query $P(Y=1|do(\textbf{X}=1))$ on Models 1, 8, & 3’ by WERM and EM4CI. $k_{lrn}$ is the learned domain sizes of latent variables.
@@ 4.3 Results
Results on models 1-8 (Figure 2(a)) are presented in Table 3. Since these models are quite small we report the total time (learning plus inference) for EM4CI. We compare against the Plug-In estimand method (see estimand expressions in Table 1), which is guaranteed to converge to the exact answer. Therefore, we expect Plug-In to produce fairly accurate results on these small models if given enough samples. We observe that the accuracy of both methods are similar at both 100 and 1,000 samples, with EM4CI being more accurate on some cases, and Plug-In on others. EM4CI had better time performance for 100 samples, but for 1,000 samples the Plug-In was faster since learning time of EM4CI was longer.
@@
The results for Models 1, 8, and 3’ 222model 3’ is the same as Model 3 in Figure 2(a), but with edge $Y\rightarrow Z$ reversed to match the hard-coded model in WERM. comparing WERM [13] to EM4CI are given in Table 4. We use domain size $d=2$, and 1,000 and 10,000 samples. Again, EM4CI produced more accurate results in several instances, though the disparities are smaller than with the Plug-In method. EM4CI was faster than WERM with 1,000 samples but slower with 10,000 samples.
Unfortunately, the available code for WERM is specific to these models, so we were unable to compare against it more generally. This also highlights a general lack of available estimand-based implementations applicable to general settings.
Figure 4: Comparing the accuracy of EM4CI and Plug-In. While both methods improve with more samples (solid to dashed lines), the error ($mad$) of EM4CI is smaller, even when compared to Plug-In with more samples.
@@
Learning: time $=199.2s$ $k_{lrn}=6$ time$=1833.7s$ $k_{lrn}=6$
@@
In Tables 5(b) and 6(b) we show results on larger models of chains, diamonds, and cone-cloud graphs. The first two tables compare EM4CI to the plug-in method for a single query over all the models. Specifically, Table 5(a) presents time and accuracy results for 1,000 samples. We see that EM4CI was consistently more accurate, and in many cases significantly better (e.g., in 45-cone-cloud). We see the same trend for 10,000 samples in Table 5(b). The superior accuracy of learning for model completion compared with the Plug-In is visualized well in Figure 4, which shows the accuracy trends for each class as a function of model sizes, for both 1,000 and 10,000 samples. We can see that generally, EM4CI using only 1,000 samples is even more accurate than Plug-In with 10,000 samples.
Again, we expect this improvement is due to the variance reduction of the estimation process. By its nature, model completion exploits more information from the causal graph than is apparent in the estimand expression alone, implicitly capturing the known conditional independence structure in its estimates of $P(\boldsymbol{V})$. We can also include simple and easy-to-impose complexity control, in the form of the latent domain sizes, to further reduce variance. In contrast, it is difficult to impose meaningful regularity or variance reduction on the plug-in estimates of $P(\boldsymbol{V})$ without creating significant computational burdens. Thus our results suggest that whenever the causal graph’s treewidth is bounded and the estimand expression has large scope functions, we should prefer using model completion. Other settings may require more study to determine the best approach.
Focusing on time, we see that the time of EM4CI is significantly more costly than Plug-In, with EM learning being the most time consuming part. Interestingly, while time grows with model size for both schemes, the inference time component remains efficient, likely due to the low treewidth of some of the models (e.g., the chains and the diamonds). In the 45-cone case, inference time is impacted more by model size, since its treewidth increases with the square root of the number of variables. Generally, for a single query, we find Plug-In has better time performance, and its time increases at a slower pace. In both methods, time also increases with sample size, e.g., when moving from 1,000 samples table to 10,000 samples.
@@
Table 7(b), compares EM4CI against Plug-In on a single query for all 4 networks. We observe the same pattern of performance as in the synthetic networks: EM4CI provides far more accurate results on all these models but with some time cost, attributed to learning the models. Evaluation for multiple queries for EM4CI and Plug-In are given in Table 8 for the $\mathbf{A}$ network. We find the same latent domain size ($k_{lrn}=4$) is selected for both sample sizes. As before, learning time grows with sample size. Accuracy results are excellent and improve with increased sample sizes as well. Note that increased sample size improves accuracy but does not impact inference time, although more data may allow selection of a more complex model (larger latent domain size), which may also improve inference accuracy but may require more time. Results for Alarm, Win95, and Barley networks are similar (see Appendix).
@@
Tables 6(b) and 8(b) highlight how learning time of EM4CI can be amortized effectively over multiple queries. The table presents EM4CI’s learning and inference time separately on the multiple queries. In contrast, the Plug-In method requires estimation of each new query from scratch, even if on the same model; see Table 6(a) and 8(a). Thus, for multiple queries, EM4CI may take far less time per query, while providing superior quality estimates. (See additional results on multiple queries in the Appendix.)
@@
Our experiments illustrate the strength of the direct-learning approach as a viable alternative for answering causal effect queries. We saw that model-completion by learning implemented in our EM4CI yields highly accurate estimates of causal effect queries on a variety of benchmarks, both synthetic networks and real life networks, on both small and large models. Moreover EM4CI shows clear superiority to the Plug-In estimands approach. In terms of time efficiency however, EM4CI was consistently slower due to learning time overhead. Yet, when answering multiple queries is desired the time overhead can be amortized over multiple queries.
@@
$|\boldsymbol{V}|=46;~{}|\boldsymbol{U}|=8;~{}d=2;~{}k=2$ treewidth $\approx 16$
@@
Learning $time=71(s)$ $k_{lrn}=4$ $time(s)=541$ $k_{lrn}=4$

### v2

@@ 1.  Initialize: $BIC_{\mathcal{B}}\leftarrow\infty$
  3.    Initialize: $LL_{\mathcal{B}^{\prime}}\leftarrow 0$
  4.    for $i=0\ldots 9$ do      (optimize LL of EM)
  5.       Initialize: $\theta_{i}$ to random parameters of $\mathcal{G}$
  6.       $(LL_{\mathcal{B}_{i}},\mathcal{B}_{i})\leftarrow\mathrm{EM}(\mathcal{G},\mathcal{D},k,\theta_{i})$
  7.       if $LL_{\mathcal{B}_{i}}>LL_{\mathcal{B}^{\prime}}$
  8.         $LL_{\mathcal{B}^{\prime}}\leftarrow LL_{\mathcal{B}_{i}},\mathcal{B}^{\prime}\leftarrow\mathcal{B}_{i}$
  9.    end for
@@ 12.          $\mathcal{B}\leftarrow\mathcal{B}^{\prime}$, $BIC_{\mathcal{B}}\leftarrow BIC_{\mathcal{B}^{\prime}}$
  15:  $\mathcal{B}_{{\boldsymbol{X}=\boldsymbol{x}}}$leftarrow truncated CBN from $\mathcal{B}$.
  16:  return $P_{\mathcal{B}_{\boldsymbol{X}=\boldsymbol{x}}}({\boldsymbol{Y}})$ computed by any PGM algorithm
@@
Since we use EM for learning we call our algorithm EM for Causal Inference (EM4CI), described in Algorithm 1. Its input is a causal diagram $\mathcal{G}$, samples $\mathcal{D}$ from $P(\boldsymbol{V})$, and an identifiable query $Q=P({\boldsymbol{Y}}\mid do({\boldsymbol{X}=\boldsymbol{x}}))$. Steps 2-14 provide the iterative learning scheme. Given samples $\mathcal{D}$ from $P(\boldsymbol{V})$, we use EM to learn a CBN $\mathcal{B}$ over the graph $\mathcal{G}$ with latent domain size $k$. EM outputs the parameters for a network denoted $\mathcal{B}_{i}$, and its corresponding log-likelihood, $LL_{\mathcal{B}_{i}}$. Since EM’s performance is known to be sensitive to initialization, we perform ten runs of EM starting from randomly generated initial parameters, and retain the model, ${\mathcal{B}^{\prime}}$, with the highest likelihood, $LL_{\mathcal{B}^{\prime}}$, as the candidate for the current latent domain size $k$ (steps 4-9), with which we calculate the $BIC_{\mathcal{B}^{\prime}}$ score (step 10). If the $BIC$ score decreases (steps 11-13), we adopt the new candidate network as $\mathcal{B}$, and the process continues with increased $k$; otherwise the current $\mathcal{B}$ is selected as the final learned network. Step 15 generates the truncated CBN $\mathcal{B}_{\boldsymbol{X}=\boldsymbol{x}}$, and step 16 employs a standard PGM inference algorithm, like join-tree decomposition [6, 9], to the CBN $\mathcal{B}_{\boldsymbol{X}=\boldsymbol{x}}$ to compute $P({\boldsymbol{Y}})$.
@@
The complexity of EM (step 6) depends on the sample size $|D|$, the variables’ domain sizes, and the number of iterations needed for convergence which is hard to predict, but complexity-wise provides only a constant factor and can be ignored. Steps 2-14 repeat the EM algorithm, which again adds only a constant factor. The most extensive computation in each iteration of EM is the Expectation step, which requires probabilistic inference and can be accomplished in time and memory exponential in the treewidth of the graph. Otherwise, approximation algorithms such as belief propagation may be used. In our experiments, exact inference was always possible. Thus, each iteration of EM takes $O(|D|\cdot nl^{w})$, where $n=|V\cup U|$ is the number of variables, $l$ bounds the variable domain sizes, and $w$ is the treewidth. Thus for $T$ iterations we have $O(T\cdot|D|\cdot n\cdot l^{w})$. Step 16 of probabilistic inference is $O(n\cdot l^{w})$ with exact inference. In summary, the complexity of algorithm EM4CI is $O(T\cdot|D|\cdot n\cdot l^{w})$. Thus, the algorithm is most effective for models with bounded treewidth. Note that the cost of the EM learning process can be amortized over multiple queries. In summary,
Theorem 2 (complexity of EM4CI).
The complexity of algorithm EM4CI is $O(T\cdot m\cdot n\cdot l^{w})$, where $m$ is the number of samples, $T$ is the number of iterations, $k$ bounds the domain size, $n$ is the total number of variables, and $w$ is the treewidth of the causal diagram.
@@
Table 1: Estimand Expressions for Models 1-8 in Figure 5 & Table 3.
@@ 4 Empirical Evaluation
We perform an extensive empirical evaluation of EM4CI and compare against estimand-based approaches such as the brute-force empirical plug-in method and a state-of-the-art scheme called weighted empirical risk minimization (WERM) [14]. All experiments were run on a 2.66 GHz processor with 8 GB of memory. All experimental code can be found here [26].
@@
Figure 3: Larger synthetic models.
@@
We generate each triplet benchmark instance by first choosing a causal diagram and a query. Then, picking domain sizes of observed and latent variables, we generate the conditional probability tables (CPTs), one per variable and its parents, to yield the full CBN. From this full model we generate samples by forward sampling [6] and then discard the values of the latent variables, yielding samples from the observed distribution. We compute the exact answer to the target query on the full CBN via an exact algorithm (e.g., join-tree [6, 9]).
The causal diagrams were selected in two ways. First, we use a set of 8 small diagrams from the causality literature [25, 15]; see Figure 5 and the Supplemental [27]. Second, we examine three scalable classes of graphs whose treewidths vary but can be controlled. These are: chain networks (treewidth 3; Figure 3(c)), diamond networks (treewidth 5; Figure 3(a)), and cone-cloud networks (treewidth $O(\sqrt{n})$, Figure 3(b)), abbreviated CH, D, CC respectively.
In the small diagrams (Figure 5), we set the domain size of the observed variables to $d=2$, and of the latent variables to $k=10$. For the chain, diamond, and cone-cloud models, we use $(d,k)=(4,10)$. The parameters of each CPT were generated by sampling from a Dirichlet distribution [6] with parameters $\alpha=[\alpha_{1},\ldots,\alpha_{n}]$, where $\alpha_{i}\in[0,1)$ selected uniformly at random. The aim is to generate conditional probabilities near the edges of the simplex (i.e., far from uniform).
@@
We also experimented with four networks created for real-world domains. The “Alarm” network was developed for on-line monitoring of patients in intensive care units [2]; the “A” network, from the UAI literature, is synthetic but known to be daunting to exact algorithms [19]; the “Barley” network was built for a decision support system for growing barley without pesticides [20]; and the “Win95” model is an expert system for printer troubleshooting in Windows 95. Since no variables in these networks are specified as latent, we designated source vertices with at least two children as latent confounding variables. As in the synthetic networks case, we generated data by forward sampling, subsequently discarding the latent variables’ values.
@@
We hand selected multiple identifiable queries per model, aiming for non-trivial ones and prioritizing those that appear to be complex. We provide the queries and their corresponding estimands in the Supplemental [27].
@@
We used the ID algorithm in the causal effect package [33] to compute the estimands. Subsequently, the plug-in method evaluates the estimand from the observational data, by plugging in the empirical conditional probabilities. Our plug-in implementation uses sparse table representations in Python, to ensure that the representation size of each estimated probability term is linear in the data size, rather than exponential in the number of variables. We also compare against a state-of-the-art scheme called WERM [14] using the author-provided implementation in R. A summary of the algorithms and implementations is given in Table 2.
@@
We report the results of EM4CI along the two phases of the algorithm; the learning phase (steps 2-14) done via EM and the inference phase (steps 15-16) done by the join tree algorithm. For the learning process, we report the selected latent domain sizes and the total time at termination. For the inference phase we report the mean absolute deviation (mad) between the true answer and the estimated answer. The measure $mad$ for a query $P({Y}\mid do({\boldsymbol{X}}))$ is computed by averaging the absolute error over all instantiations of the intervened and queried variables, $P({Y=y}\mid do({\boldsymbol{X}}=\boldsymbol{x}))$ for $\boldsymbol{x},y\in\mathcal{D}({\boldsymbol{X}})\times\mathcal{D}(Y)$. We also report inference time for each query. We incremented $k$ by steps of two ($k=2,4,6,8,10\ldots)$ in the EM4CI algorithm.
@@
1 2 0.0037 0.48 0.0104 1.9 2 0.0032 3.1 0.0025 2.3
2 2 0.1832 1.86 0.1436 2.3 2 0.0490 8.4 0.0867 2.0
3 2 0.1288 0.93 0.0569 1.1 2 0.0040 3.6 0.0039 0.7
4 2 0.1819 1.82 0.1469 2.3 2 0.1438 12.0 0.0704 2.1
5 2 0.4910 1.65 0.5000 2.0 2 0.0044 17.3 0.0058 2.2
6 2 0.2663 0.30 0.3930 2.0 2 0.1263 0.5 0.1319 2.1
7 2 0.2520 0.78 0.2509 1.9 2 0.0891 7.1 0.0238 2.0
8 2 0.1372 0.63 0.1579 2.0 2 0.2340 4.7 0.1303 1.9
Table 4: Results of absolute error on Query $P(Y=1|do(\textbf{X}=1))$ on Models 1, 8, & 3’ by WERM & EM4CI. $k_{lrn}$ is the learned domain sizes of latent variables.
@@ 4.3 Results
Results on models 1-8 (Figure 5) are presented in Table 3. Since these models are quite small we report the total time (learning plus inference) for EM4CI. We compare against the Plug-In estimand method (see estimand expressions in Table 1), which is guaranteed to converge to the exact answer. Therefore, we expect Plug-In to produce fairly accurate results on these small models if given enough samples. We observe that the accuracy of both methods are similar at both 100 and 1,000 samples, with EM4CI being more accurate on some cases, and Plug-In on others. EM4CI had better time performance for 100 samples, but for 1,000 samples the Plug-In due to the learning time of EM4CI.
@@
The results for Models 1, 8, and 3’ 111model 3’ is the same as Model 3 in Figure 5, but with edge $Y\rightarrow Z$ reversed to match the hard-coded model in WERM. comparing WERM [14] to EM4CI are given in Table 4. We use domain size $d=2$, and 1,000 and 10,000 samples. Again, EM4CI produced more accurate results in several instances, though the disparities are smaller than with the Plug-In method. EM4CI was faster than WERM with 1,000 samples but slower with 10,000 samples. Unfortunately, the available code for WERM is specific to these models, so we were unable to compare against it more generally.
Error (mad)
Refer to captionNumber of Variables Refer to captionRefer to captionRefer to caption
Figure 4: Comparing the accuracy trends of EM4CI and Plug-In as a function of the model size. While both methods improve with more samples (solid to dashed lines), the error ($mad$) of EM4CI is smaller, even when compared to Plug-In with more samples.
@@
Learning: time(s) $=199$ $k_{lrn}=6$ time(s)$=1833$ $k_{lrn}=6$
@@
In Tables 5(b) and 6(b) we show results on larger models of chains, diamonds, and cone-cloud graphs. The first two tables compare EM4CI to the plug-in method for a single query over all the models. Specifically, Table 5(a) presents time and accuracy results for 1,000 samples. We see that EM4CI was consistently more accurate, and in many cases significantly better (e.g., in 45-cone-cloud). Table 5(b) shows the same trend for 10,000 samples. The superior accuracy of learning for model completion over Plug-In is clearly visible in Figure 4, which shows the accuracy trends for each class as a function of model sizes, for both 1,000 and 10,000 samples. We see that generally, EM4CI using only 1,000 samples is even more accurate than Plug-In with 10,000 samples.
Again, we expect that this improvement is due to the variance reduction of the estimation process. By its nature, model completion exploits more information from the causal graph than is apparent in the estimand expression alone, implicitly capturing the known conditional independence structure in its estimates of $P(\boldsymbol{V})$. We can also include simple and easy-to-impose complexity control, in the form of the latent domain sizes, to further reduce variance. In contrast, it is difficult to impose meaningful regularity or variance reduction on the plug-in estimates of $P(\boldsymbol{V})$ without creating significant computational burdens. Taking it all together, our results suggest that whenever the causal graph’s treewidth is bounded and the estimand expression has large scope functions, we should prefer using model completion. Other settings may require more study to determine the best approach.
Focusing on time, we see that the time of EM4CI is significantly larger than Plug-In, with EM learning being the most time consuming part. Interestingly, while time grows with model size for both schemes, inference time remains efficient, likely due to the low treewidth of some of the models (e.g., the chains and the diamonds). In the 45-cone case, inference is impacted more by model size, since its treewidth increases with the square root of the number of variables. Generally, for a single query, we find that the Plug-In is superior time-wise, and its time increases at a slower pace with model size. Finally, in both methods, time increases with sample size as expected.
@@
Table 7(b) compares EM4CI against Plug-In on a single query for all 4 networks. We observe the same pattern of performance as in the synthetic networks: EM4CI provides far more accurate results on all real-world models but with some time cost, attributed to learning. Evaluation for multiple queries are given in Table 8 for the $\mathbf{A}$ network. We find that the same latent domain size ($k_{lrn}=4$) is selected for both sample sizes. As before, learning time grows with sample size and accuracy results are excellent and are improving with increased sample sizes. Note, however, that increased sample size does not impact inference time. Results for Alarm, Win95, and Barley networks are similar and can be found in the supplemental [27].
@@
Tables 6(b) and 8(b) highlight how learning time can be amortized effectively over multiple queries. The table presents EM4CI’s learning and inference time separately on the multiple queries. In contrast, the Plug-In method requires estimation of each new query from scratch, even if on the same model; see Table 6(a) and 8(a). Thus, for multiple queries, EM4CI may take far less time per query, while providing superior quality estimates. (See additional results on multiple queries in the Supplemental [27].)
@@
Our experiments illustrate that model-completion by learning implemented in EM4CI yields highly accurate estimates of causal effect queries on a variety of benchmarks, including both synthetic and real-world networks of varying sizes. Moreover EM4CI shows clear superiority to the Plug-In estimands approach. In terms of time efficiency however, EM4CI was consistently slower due to learning time overhead. Yet, when answering multiple queries is desired the time overhead can be amortized over multiple queries.
@@
$|\boldsymbol{V}|=46;~{}|\boldsymbol{U}|=8;~{}d=2;~{}k=2$, treewidth $\approx 16$
@@
Learning $time(s)=71$ $k_{lrn}=4$ $time(s)=541$ $k_{lrn}=4$